# Quantum latent distributions in deep generative models

### v1

@@ IV Benchmarking GANs with quantum latent distributions
Figure 3: Left: We compare GANs that use different latent space distributions on different datasets. We focus on quantum latent distributions generated by boson samplers, a type of quantum processor. Right: in a boson sampler, several identical photons enter an optical interference network and a measurement is performed to determine where the photons leave the network. Classical simulations of this probabilistic process are inefficient, and intractable beyond a few tens of photons.
In this section, we empirically investigate whether the use of a quantum latent distribution can lead to a performance improvement over relevant classical distributions. We consider generative adversarial networks (GANs), where the use of different latent distributions is well-studied. We perform two experiments: one on a small-scale toy dataset produced by a quantum process, and one on the QM9 quantum chemistry dataset. For both experiments, as illustrated in Figure 3 we compare different types of latent space distributions that produce samples $z$ of length $L$ ($z_{dim}=L$). We select three classical latent distributions and one quantum distribution:
@@
  • Bernoulli: random uniform bit strings (discrete), $z\in\{0,1\}^{L}$
  • Dist. sampler: arising from distinguishable photons sent into a size-$L$ interferometer
  • Boson sampler: arising from indistinguishable photons sent into a size-$L$ interferometer
We select the Gaussian distribution since this is the most commonly used classical latent distribution. The other two distributions represent two intermediate steps that bring us closer to the quantum distribution. The Bernoulli distribution is a discrete rather than continuous latent distribution, which yielded good performance in [23]. The distribution arising from sending distinguishable single photons, which do not interfere with each other (for example photons with different frequencies), into a size-$L$ interferometer is a discrete distribution that can be produced efficiently classically. The difference between the distinguishable and indistinguishable photon distributions is due to quantum interference between photons when they are indistinguishable, such that a difference in performance between these two latent distributions can be attributed to the different statistics arising from this interference.
To ensure an apples-to-apples comparison, for all experiments our models differ only through the nature of the process that produced the samples that are sent to the generator. Moreover, both types of photonic distributions arise from randomly initialized interference circuits. To ensure that performance differences arise from general properties of these distributions and not from specific circuit realizations, these circuits are re-sampled for every seed (with an exception for the largest-sized latent spaces, see appendix for details). We note that these circuits could additionally be trained [43], although training methods for quantum circuits often scale poorly [44]. We do not train them in this work since this could lead to an unfair advantage in benchmarking against the static Gaussian and Bernoulli distributions.
IV.1 Experiment on a toy dataset
For our small-scale experiment, we consider two toy datasets: a "Quantum dataset" produced by simulating [45] 8 identical photons interfering in a 16-channel random optical circuit, and a "Bernoulli dataset" produced by a 16-dimensional Bernoulli distribution. The optical circuits used for the latent and data distributions are independently sampled, preventing trivial identity mapping. Since the two target distributions are discrete, to quantify the performance of a trained model we use the average distance between the outputs of the generator and their nearest integers. A smaller distance indicates that the generator has been more successful in learning the discrete nature of the distribution.
@@ IV.1 Experiment on a toy dataset
Gaussian Bernoulli Dist. sampler Boson sampler
@@
Our results comparing the performance of the trained models are shown in Table 1. A first observation is that the classical dataset is easier to learn, with smaller distances and smaller differences between the latent distributions. On the harder quantum dataset, we observe significant differences between the latent distributions, with the quantum latent distribution achieving the best results. In particular, the quantum latent distribution outperforms the latent distribution that uses distinguishable photons. This implies that the quantum interference between photons, which is the only difference between these two distributions, is a useful resource that the model used to achieve better performance on the quantum dataset. These results validate the expectation that a quantum distribution can be a more suitable latent distribution when the data also arises from a quantum process.
@@ IV.2 Experiments on the QM9 dataset
Next, we consider the QM9 quantum chemistry dataset [46]. Generating chemical structures with specific properties is a challenging problem, partly because the underlying dynamics are described by quantum mechanical principles. As such, quantum chemistry is a promising field for realizing the potential advantages of a quantum latent distribution. The QM9 dataset benefits from being a reasonably small-size dataset that is amenable to extensive benchmarking, and which has been previously studied within the context of GANs [47, 48, 49]. We use a GAN architecture based on MolGAN [47], with some improvements described in the appendix. The generator is a feedforward neural network which is loosely based on the requirements of theorem 1: all layers except the final layer are non-decreasing, and we use a LeakyReLU activation function. The discriminator is a relational graph convolutional network.
@@ IV.2.1 Type and size of the latent distribution
We first compare results over different latent distributions and different sizes: 16, 32 and 48. For the distinguishable and indistinguishable photon distributions, we consider unstructured optical circuits, with transfer matrices described by Haar-random unitary matrices, and with half as many input photons as optical channels. For example, the size-16 latent is produced by 8 photons in 16 channels. We use the algorithm in [45] to simulate these systems.
@@
Latent Type $z_{dim}$ FCD $\downarrow$ # Valid & unique $\uparrow$ # Novel $\uparrow$
Boson samp. 16 1.160 ± 0.06 2522 ± 65 1331 ± 37
Dist. samp. 16 1.333 ± 0.07 1954 ± 103 1067 ± 54
@@
Boson samp. 32 1.536 ± 0.08 1791 ± 106 951 ± 37
Dist. samp. 32 1.533 ± 0.06 1633 ± 81 919 ± 45
@@
Boson samp. 48 1.696 ± 0.08 1528 ± 65 856 ± 40
Dist. samp. 48 1.713 ± 0.06 1307 ± 77 746 ± 43
@@
Our results are shown in table 2, with additional results such as training curves in the appendix. We find that, for all 3 latent space sizes, the quantum distribution outperforms the classical distributions. The largest performance improvement occurs for the size-16 latent space, where the quantum distribution outperforms all other latent distributions on all metrics. Though the gap narrows for larger latent spaces, and overall performance decreases, the size-48 quantum latent distribution still leads to a higher number of valid and unique generated molecules than the classical baselines. Moreover, as shown by the improvement over distinguishable photons, the statistics arising from quantum interference is a factor that improves performance. These results show that a quantum latent distribution can outperform other relevant classical latent distributions on a chemistry data set.
These experiments were limited to size-48 latents due to the exponentially increasing complexity of simulating a quantum processor. Generating 1 million size-48 samples with the algorithm in [45] took about 3.75 hours on 100 AMD EPYC™ 7003 processors, which may be close to a regime where a near-term quantum processor would run faster. The persistence of a small advantage at size 48 on the number of valid, unique and valid molecules suggests that an advantage over the classical distributions may still be achievable for a quantum processor of a size that is not practically simulable. Though in this case a smaller latent space yields improved results, for other models and datasets larger-scale latents have been shown to work best [23], which would motivate further exploration of this approach for larger-scale quantum distributions.
@@ IV.2.2 Type of quantum circuit
To understand whether the results presented in the previous section are specific to Haar-random optical circuits or are due to more general properties of the statistics of photon interference, we also ran experiments with different types of optical circuits. We simulate two realizations of a boson sampling circuit based on optical delay lines, which have been proposed and demonstrated as an experimentally feasible route to large-scale boson sampling [14, 51, 52]. These realizations respectively have two delay lines of the same length (a "1-1" configuration), and three delay lines in a "1-3-9" configuration where each line delays each photon by three times the previous line.
Table 3: Metrics (± standard error over 20 seeds) for different latents using different quantum circuits
Latent Type $z_{dim}$ FCD $\downarrow$ # Valid & unique $\uparrow$ # Novel $\uparrow$
Boson samp. 1-1 16 1.245 ± 0.06 1375 ± 72 692 ± 34
Dist. samp. 1-1 16 1.291 ± 0.07 1182 ± 96 580 ± 47
Boson samp. 1-3-9 16 1.187 ± 0.06 1990 ± 106 1063 ± 58
Dist. samp. 1-3-9 16 1.376 ± 0.09 1459 ± 100 775 ± 53
Our results are shown in Table 3. We find that for each type of optical circuit, the quantum distribution outperforms the classical distribution arising from distinguishable photons. This shows that for this model and dataset the statistics arising from quantum interference confer an advantage that does not depend on the specifics of the circuit architecture. We also find that the circuit architecture has an impact on the overall performance of the latent distribution, with the Haar-random circuit performing best overall for size-16 latents. This suggests that developing methods for circuit architecture search or joint training of the circuit with the neural network could lead to improved results.
IV.2.3 Experiments with a real quantum processor
Realistic boson sampling systems are subject to a range of imperfections such as optical loss, limited photon number resolution, and imperfect indistinguishability. To assess the extent to which our previous results are transferrable to a real quantum processor, we performed experiments on a real boson sampling system, a commercially available ORCA Computing PT-2. This system can interfere 16 photons in 32 channels and its optical network consists of optical delay lines in the "1-1" configuration. More information about this system can be found in the appendix. We collected half a million samples from this system with a single circuit realization, and compared the results to comparable size-32 latent distributions.
@@ IV.2.3 Experiments with a real quantum processor
Latent Type $z_{dim}$ FCD $\downarrow$ # Valid & unique $\uparrow$ # Novel $\uparrow$
Boson samp. 1-1 (real) 32 1.410 ± 0.07 1962 ± 75 1084 ± 45
Boson samp. 1-1 (simulated) 32 1.414 ± 0.07 1903 ± 66 1073 ± 36
Dist. samp. 1-1 32 1.560 ± 0.09 1638 ± 102 900 ± 50
@@
Our results are shown in Table 4. We find that the results from the real quantum processor closely match the simulated quantum processor despite experimental imperfections. Both quantum distributions also outperform the classical distributions. These results indicate that a performance advantage achieved with a simulated quantum system can provide a good indication of the performance of a real quantum processor, which can be helpful for extrapolating the performance of small-scale simulable distributions to larger-scale, non-simulable distributions.
V Quantum distributions for diffusion and flow matching
In contrast to GANs where the latent distribution can be straightforwardly swapped out for a different distribution, the specifics of how such a process can be carried out for other models are less clear. This has restricted most studies of quantum latent distributions in deep generative models only to GANs. Here, we broaden the set of models that can make use of this method. Though the experiments in this section do not exhibit a performance advantage with a quantum latent distribution, they demonstrate the feasibility of this approach and motivate further exploration.
V.1 Denoising Diffusion GANs
We first investigate diffusion models, which often achieve superior performance to GANs in fields such as images [53] and protein structure generation [54]. Latent diffusion models generally make Gaussianity assumptions on the latent distribution [3], such that the use of a quantum distribution is not straightforward. However, it has been observed that the reverse process, when considered over large time steps, involves modeling a complex non-Gaussian distribution, for which a GAN can be well suited [55]. We thus focus on Denoising Diffusion GANs (DDGANs) [55], which use the standard forward diffusion process but use a GAN to perform the reverse process. They achieve results comparable to standard diffusion models with faster inference speeds, with a model that is easier to train than other GANs. This work aligns with the hypothesis [10] that at least part of the observed improvement in diffusion models over GANs may stem from the additional computation at inference due to the recursive sampling process rather than a fundamental paradigm difference.
We investigated unconditional generation on the CIFAR10 dataset [56] for different latent distributions, using the model architecture in [55]. Where the original model used a size-256 latent space, our experiments used a size-48 latent space due to the classical complexity of simulating larger quantum circuits. We found results comparable to those in [55]. This experiment shows the feasibility of this approach and motivates future work to identify datasets and metrics where a quantum latent distribution can improve results.
V.2 Flow matching
We also identify flow matching models as being well suited to use a quantum distribution in the latent space. Flow matching has emerged as a flexible tool for producing high-quality data with potentially fewer inference steps than diffusion models. Though many such models require a source distribution that is Gaussian, recently flow matching models that can use an arbitrary source distribution have been proposed [5, 6]. To illustrate the feasibility of a quantum latent distribution as a source distribution, Figure 4 represents a small-scale model trained using the method from [27] to map 2D-projected quantum latent vectors to the "moons" dataset. Further work would be required to identify models and datasets on which a quantum distribution can provide a benefit.
VI Discussion
Here, we discuss two potential limitations of using a quantum latent distribution for a generative model. These are the extent to which other classical distributions could be used instead of the quantum distribution to achieve improved performance, and whether the quantum computer’s output could be classically simulated to achieve the same result.
First, though we compared the quantum distribution to a range of progressively more similar classical distributions, we did not perform an exhaustive search of different classical latent distributions. More complex classical distributions could be used, for example trained distributions using an autoencoder at the expense of additional training [10]. Some of these methods may yield improved performance on this problem. However, investigating a wide range of alternative classical distributions is impractical and computationally expensive, while the increasing availability of quantum computers [57] means that they are progressively becoming a practical alternative. Furthermore, this work provides a principled approach for identifying the settings that are most likely to benefit from a quantum distribution, namely datasets originating from a physical process that is poorly described by classical mechanics and highly multimodal datasets. As such, in these settings the use of a quantum computer may be a more effective use of resources than investigating a wider range of classical distributions.
@@ VI Discussion
Additionally, though our comparison between distinguishable and indistinguishable photons allows us to show that the statistics arising from quantum interference can improve performance, other classical distributions may be able to reproduce similar statistics without a quantum processor. In particular, classical methods for simulating photonic quantum processors in the presence of noise have been developed [58, 59]. These methods may be able to achieve a similar performance improvement, however for producing latent distributions they are likely to be impractical. The approach of [58] reproduces the low-order statistics of a boson sampling distribution with a quadratic overhead compared to a linear overhead at most with a real physical system. The alternative approach of [59] required 9 minutes on 288 GPUs to set up the simulation initially. Neither approach is thus likely to run faster than a real quantum processor. Though the speed of generating samples is less of a concern for the apples-to-apples benchmarking setting that we selected in this work where the latent distribution is not trained, this can be a major limitation for more advanced variational settings. Therefore, if the latent distribution were to be trained jointly with the neural network, for example using the method in [43], these existing classical approximation methods would not be suitable.
VII Conclusion
In this work, we examined the use of quantum distributions in the latent space of deep generative models, providing both theoretical and empirical arguments for a potential benefit. We showed that under some assumptions on the generator architecture, a quantum latent distribution can allow a model to produce a broader class of data distributions. Through benchmarks on several datasets, we empirically found that quantum latent distributions can indeed outperform relevant classical distributions.
Though these results establish that near-term quantum processors can help expand the capabilities of generative models, several questions remain to be addressed and could be the focus of future work. First, though we propose two mechanisms through which a quantum latent distribution could be used to achieve a practical performance improvement, a deeper theoretical understanding of the inductive bias that these distributions provide and how the non-factorizability of the latent distribution affects the outputs would be beneficial. Second, we anticipate that, in the same way that trained classical latent distributions can lead to improved results [10], developing methods to train the quantum processor jointly with the generator would also yield better performance. Finally, extending our experiments on diffusion and flow matching to larger-scale models and different data modalities would help evaluate the potential of this approach to improve the current state of the art on challenging real-world problems.
VIII Author contributions and acknowledgements

### v2

@@ IV Benchmarking GANs with quantum latent distributions
In this section, we empirically investigate whether the use of a quantum latent distribution can lead to a performance improvement over relevant classical distributions. We consider generative adversarial networks (GANs), where the use of different latent distributions is well-studied. Though GANs are no longer state of the art for several types of data, they provide a straightforward mapping from latent distribution to data distribution that facilitates apples-to-apples comparisons between latent distributions.
We perform two experiments: one on a small-scale synthetic dataset produced by a quantum process, and one on the QM9 quantum chemistry dataset. Based on the intuition developed in the previous section, these datasets have properties that make them a good fit for testing a quantum latent distribution. Both datasets arise from fundamentally quantum mechanical processes and are highly non-uniform.
For both experiments, we compare different types of latent distributions that produce samples $z$ of length $d_{z}=L$. We select one quantum distribution and three classical latent distributions:
  • Quantum: arising from indistinguishable photons sent into a size-$L$ interferometer
  • Photonic: arising from distinguishable photons sent into a size-$L$ interferometer
  • Bernoulli: random uniform bit strings (discrete), $z\in\{0,1\}^{L}$
@@
The distribution arising from sending distinguishable single photons, which do not interfere with each other (for example photons with different frequencies), into an interferometer is a discrete distribution that can be produced efficiently classically. The only difference between the distinguishable and indistinguishable photon distributions is quantum interference between photons when they are indistinguishable. As such, a difference in performance between the photonic and quantum latent distributions can be entirely attributed to the different statistics arising from quantum interference.
To ensure an apples-to-apples comparison, for all experiments our models differ only through the nature of the latent distribution. Moreover, both the quantum and photonic distributions arise from randomly initialized interference circuits. To ensure that performance differences arise from general properties of these distributions and not from specific circuit realizations, these circuits are re-sampled for every seed (with an exception for the largest-sized latent spaces, see appendix for details). We note that these circuits could additionally be trained [20], although training methods for quantum circuits often scale poorly [40]. We do not train them in this work since this could lead to an unfair advantage in benchmarking against the static Gaussian and Bernoulli distributions.
IV.1 Experiment on synthetic datasets
For our small-scale experiment, we consider two synthetic datasets: a "Quantum dataset" produced by simulating [13] 8 identical photons interfering in a 16-channel random optical circuit, and a "Bernoulli dataset" produced by a 16-dimensional Bernoulli distribution. The optical circuits used for the latent and data distributions are independently sampled, preventing a trivial identity mapping from the latent space to the data space. Since the two target distributions are discrete, to quantify the performance of a trained model we use the average distance between the outputs of the generator and their nearest integers. A smaller distance indicates that the generator has been more successful in learning the discrete nature of the distribution.
@@ IV.1 Experiment on synthetic datasets
Gaussian Bernoulli Photonic Quantum
@@
Our results comparing the performance of the trained models are shown in Table 1. A first observation is that the classical dataset is easier to learn, with smaller distances and smaller differences between the latent distributions. On the harder quantum dataset, we observe significant differences between the latent distributions, with the quantum latent distribution achieving the best results. We find that the quantum latent distribution outperforms the latent distribution that uses distinguishable photons. This implies that the statistics arising from quantum interference between photons are a useful resource that the model used to achieve better performance on the quantum dataset. These results validate the expectation that a quantum distribution can be a suitable latent distribution when the data also arises from a quantum process.
@@ IV.2 Experiments on the QM9 dataset
Next, we consider the QM9 quantum chemistry dataset [49]. Generating chemical structures with specific properties is a challenging problem, partly because the underlying dynamics are described by quantum mechanical principles. The QM9 dataset is a reasonably small-size dataset that is amenable to extensive benchmarking, and which has been previously studied within the context of GANs [15, 22, 33]. We use a GAN architecture based on MolGAN [15], with some improvements described in the appendix. The generator is a feedforward neural network which is based on the requirements of Theorem 1: all layers are non-decreasing, and we use a LeakyReLU activation function. The discriminator is a relational graph convolutional network.
@@ IV.2.1 Type and size of the latent distribution
We first compare results over different latent distributions and different latent space sizes: 16, 32 and 48. For the distinguishable and indistinguishable photon distributions, we consider unstructured optical circuits, with transfer matrices described by Haar-random unitary matrices, and with half as many input photons as optical channels. For example, the size-16 latent is produced by 8 photons in 16 channels. We use the algorithm in [13] to simulate these systems.
@@
Latent Type $d_{z}$ FCD $\downarrow$ # Valid & unique $\uparrow$ # Novel $\uparrow$
Quantum 16 1.160 ± 0.06 2522 ± 65 1331 ± 37
Photonic 16 1.333 ± 0.07 1954 ± 103 1067 ± 54
@@
Quantum 32 1.536 ± 0.08 1791 ± 106 951 ± 37
Photonic 32 1.533 ± 0.06 1633 ± 81 919 ± 45
@@
Quantum 48 1.696 ± 0.08 1528 ± 65 856 ± 40
Photonic 48 1.713 ± 0.06 1307 ± 77 746 ± 43
@@
Our results are shown in Table 2, with additional results such as training curves in the appendix. We find that, for all 3 latent space sizes, the quantum distribution achieves the best results. The largest performance improvement occurs for the size-16 latent space, where the quantum distribution outperforms all other latent distributions on all metrics. Though the gap narrows for larger latent spaces, and overall performance decreases, the size-48 quantum latent distribution still leads to a higher number of valid and unique generated molecules than the classical baselines. Moreover, as shown by the improvement over distinguishable photons, the statistics arising from quantum interference is a factor that improves performance. These results indicate that a quantum latent distribution can outperform other relevant classical latent distributions on chemistry data.
@@ IV.2.2 Type of quantum circuit
To understand whether the results presented in the previous section are specific to Haar-random optical circuits or are due to more general properties of the statistics of photon interference, we also run experiments with different types of optical circuits. We simulate two realizations of boson sampling circuits based on optical delay lines, which have been proposed and demonstrated as an experimentally feasible route to large-scale boson sampling [39, 16, 45]. These realizations respectively have two delay lines of the same length (a "1-1" configuration), and three delay lines in a "1-3-9" configuration where each line delays each photon by three times the previous line.
Table 3: Metrics (± standard error over 20 seeds) for different latents using different optical circuits
Latent Type $d_{z}$ FCD $\downarrow$ # Valid & unique $\uparrow$ # Novel $\uparrow$
Quantum 1-1 16 1.245 ± 0.06 1375 ± 72 692 ± 34
Photonic 1-1 16 1.291 ± 0.07 1182 ± 96 580 ± 47
Quantum 16 1.187 ± 0.06 1990 ± 106 1063 ± 58
Photonic 1-3-9 16 1.376 ± 0.09 1459 ± 100 775 ± 53
Our results are shown in Table 3. We find that for each type of optical circuit, the quantum distribution outperforms the classical distribution arising from distinguishable photons. This shows that for this model and dataset the statistics arising from quantum interference confer an advantage that does not depend on the specifics of the circuit architecture. We also find that the circuit architecture has an impact on the overall performance of the latent distribution, with the Haar-random circuit from table 2 performing best overall for size-16 latents. This suggests that developing methods for circuit architecture search or joint training of the circuit with the neural network could lead to improved results.
IV.2.3 Using a real quantum processor
Realistic boson sampling systems are subject to a range of imperfections such as optical loss, limited photon number resolution, and imperfect indistinguishability. To assess the extent to which our previous results are transferrable to a real quantum processor, we performed experiments on a commercially available boson sampling system. This system can interfere 16 photons in 32 channels and its optical network consists of optical delay lines in the "1-1" configuration. More information about this system can be found in the appendix. We collected half a million samples from this system with a single circuit realization, and compared the results to comparable size-32 latent distributions.
@@ IV.2.3 Using a real quantum processor
Latent Type $d_{z}$ FCD $\downarrow$ # Valid & unique $\uparrow$ # Novel $\uparrow$
Quantum 1-1 (real) 32 1.410 ± 0.07 1962 ± 75 1084 ± 45
Quantum 1-1 (simulated) 32 1.414 ± 0.07 1903 ± 66 1073 ± 36
Photonic 1-1 32 1.560 ± 0.09 1638 ± 102 900 ± 50
@@
As shown in Table 4, the results from the real quantum processor closely match the simulated system despite experimental imperfections. Both quantum distributions also outperform the classical distributions. These results indicate that a performance advantage achieved with a simulated quantum system can provide a good indication of the performance of a real quantum processor, which can be helpful for extrapolating the performance of small-scale simulable distributions to larger-scale, non-simulable distributions.
V Discussion
While the magnitude of the improvement varies across latent dimensionalities and experimental settings, our empirical study shows that quantum latent distributions can match or exceed the performance of relevant classical baselines on generative chemistry tasks. Notably, quantum interference effects appear to be a meaningful source of inductive bias for the QM9 dataset. These results provide experimental support for the theoretical arguments of Section 3: that quantum latent distributions may be beneficial in settings where data distributions are strongly correlated, multimodal, or rooted in quantum mechanical processes.
These experiments were limited to size-48 latents due to the exponentially increasing complexity of simulating a quantum processor. Generating 1 million size-48 samples with the algorithm in [13] took about 3.75 hours on 100 AMD EPYC™ 7003 processors, which may be close to a regime where a near-term quantum processor would run faster. The persistence of a small advantage at size 48 on the number of valid, unique and valid molecules suggests that an advantage over the classical distributions may still be achievable for a quantum processor of a size that is not practically simulable. Though for these experiments a smaller latent space yields improved results, for other models and datasets larger-scale latents have been shown to work best [10], which would motivate further exploration of this approach for larger-scale quantum distributions.
Though our work has focused on GANs, the theoretical understanding of quantum latent distributions developed in this work can also be extended to other types of models. Flow matching models in particular have been shown to be highly sensitive to the choice of latent distribution [32]. To show the feasibility of quantum latent distributions in other models, in the appendix we demonstrate implementations of diffusion and flow matching models that use a quantum latent distribution.
V.1 Classical alternatives
@@
Though our comparison between distinguishable and indistinguishable photons show that the statistics arising from quantum interference can improve performance, a relevant question is whether other classical distributions may be able to reproduce similar statistics without a quantum processor. In particular, classical methods for simulating photonic quantum processors in the presence of noise have been developed [58, 46]. These methods may be able to achieve a similar performance improvement, however they are likely to be impractical for machine learning applications. The approach of [58] reproduces the low-order statistics of a boson sampling distribution with a quadratic overhead compared to a linear overhead at most with a real physical system. The alternative approaches of [46, 18] require a significant amount of time to initially set up the simulation. Neither approach is thus likely to run faster than a real quantum processor. In a more advanced setting where the latent distribution were to be trained jointly with the neural network, for example using the method in [20], slower speeds are likely to make these existing classical approximation methods unsuitable.
An alternative to approximated quantum distributions to achieve a performance improvement would be to use entirely different classes of classical distributions. More complex classical distributions could be used, for example trained distributions using an autoencoder at the expense of additional training [25]. However, testing a wide range of alternative classical distributions is impractical and computationally expensive. On the other hand, quantum computers are progressively becoming more widely available [17, 39, 7, 53]. In the settings identified in this work in which a quantum latent distribution could be expected to yield good results, a quantum computer may be an effective use of resources.
VI Conclusion
In this work, we have developed an understanding of how quantum latent distributions can have an impact on the outputs of generative models. While recent experiments had already observed a performance advantage on some datasets, here we established theoretical foundations for this approach, provided intuition as to the possible causes of this improved performance, and presented rigorous benchmarks that validate these intuitions.
These results motivate further exploration of quantum latent distributions in modern generative architectures. More broadly, our results highlight that the latent space is an important component of inductive bias in a model, and that quantum distributions provide a distinct and practically accessible family of biases. Exploring this space in larger and more expressive generative models is a compelling direction for future work.
VII Author contributions and acknowledgements